triangle abc has sides 10,24, and 26 cm long. a rectangle that has an area equal to that of the triangle has width 3cm . find the perimeter of the rectangle
Answer by mananth(16946)

triangle abc has sides 10,24, and 26 cm long. a rectangle that has an area equal to that of the triangle has width 3cm . find the perimeter of the rectangle
s= half perimeter = 60 /2 =30
area of traiangle = sqrt(30*(20)*(6)*(4))
Area = 120sq.cm
..
Area of rectangle = L*w
120=3*L
L= 40 cm